3. The order reads Dopamine 2.5mcg/kg/min. The bag is labeled: 800mg in 500mL NS. The patient weighs 187 pounds. How many mL/hour will infuse?
Answer #1

Weight of the patient =187 Pounds

We require the weight in kg

Lets convert from pounds into kg,

1 kg = 2.2 pounds

187 pounds = ? kg

= 187/2.2

= 85 kg

Ordered dose of dopamine = 2.5 mcg/kg/min

Dopamine concentration in the bag = 800mg in 500 ml of NS

So 1 mL of NS = 800/500

= 1.6mg of Dopamine in 1 mL of NS

Drip rate of dopamine (mL/hr)= (Ordered dose in mcg/kg/min) x (weight in kg) x (60) / (Concentration in mg/mL) x (1000)

=2.5 x 85 x 60 /1.6 x 1000

=255 / 32

The drip rate for dopamine = 7.96 mL/hr

Hence, the Dopamine will be infused at the rate of 7.96mL/hr.